# publications.bib

### v3

@@ 1. Introduction
for all possible $b\in B(G,\mu)_{\mathrm{un}}$. This makes the situation much more complicated; in fact, for non-split $G$, the basic element could be unramified, and in this case the Igusa variety is just a profinite set, hence the problem of torsion vanishing for the contribution of the basic locus is completely reduced to controlling the generic part of the torsion cohomology of the local shtuka space attached to the basic element.
@@
Such control of the cohomology of shtuka spaces with torsion coefficients for these more general situations was attained in [Ham2]. In order to understand this, it is helpful to move away from the language of isotypic parts of shtuka spaces and consider the action of Hecke operators on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, the derived category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves on $\mathrm{Bun}_{G}$. Since we are interested in cohomology localized at a generic maximal ideal $\mathfrak{m}$, we construct in appendix A a full-subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ together with an idempotent localization map $(-)_{\phi_{\mathfrak{m}}}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ such that, on smooth irreducible representations, the localization map is either an isomorphism or 00 depending on if the representation has Fargues-Scholze parameter conjugate to $\phi_{\mathfrak{m}}$ or not (Lemma 4.2 (1)). We let $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ denote the full subcategory of ULA objects, where we recall by [FS, Theorem V.7.1], that this is equivalent to insisting that the restrictions to all the HN-strata indexed by $b\in B(G)$ are valued in the full subcategories $\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ of admissible complexes (i.e the invariants under all open compacts $K\subset J_{b}(\mathbb{Q}_{p})$ is a perfect complex). Using the results of [Ham2], we show, under various technical hypothesis including the genericity of $\mathfrak{m}$, that one has a direct sum decomposition:
$\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\simeq\bigoplus_{b\in B(G)_{\mathrm{un}}}\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}.$
@@
More precisely, we show that the !!! and * push-forwards with respect to the inclusion of HN-strata agree on this sub-category, and so the excision semi-orthogonal decomposition splits on $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$. This decomposition is a refinement of the fact mentioned above that only the shtuka spaces corresponding to the unramified elements $b\in B(G,\mu)_{\mathrm{un}}$ can contribute to the generic localization of the cohomology of the Shimura variety. The desired control of the shtuka spaces is now in turn encoded in understanding how Hecke operators interact with a perverse $t$-structure on $\mathrm{Bun}_{G}$ after restricting to the localized category $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$.
@@
[…]
where $\mathrm{Hck}_{G,\leq\mu}$ is the stack parametrizing modifications $\mathcal{E}_{1}\rightarrow\mathcal{E}_{2}$ of a pair of $G$-bundles with meromorphy bounded by $\mu$ at the closed Cartier divisor defined by the fixed untilt given by $C$, and $h_{\mu}^{\rightarrow}$ (resp. $h_{\mu}^{\leftarrow}$) remembers $\mathcal{E}_{1}$ (resp. $\mathcal{E}_{2}$). We define
@@
where $E_{\mu}$ denotes the reflex field of $\mu$ and $\mathcal{S}_{\mu}$ is the sheaf on $\mathrm{Hck}_{G,\leq\mu}$ attached to the highest weight tilting module $\mathcal{T}_{\mu}\in\mathrm{Rep}_{\overline{\mathbb{F}}_{\ell}}(\hat{G})$ of highest weight $\mu$ by geometric Satake. The action of Hecke operators commutes with the action of excursion operators and therefore the action of the spectral Bernstein center. Moreover, it preserves the subcategory of ULA objects. It follows that we have an induced map
@@
We are almost ready to state the result on Hecke operators we will need. To do this, we recall that $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ has a natural perverse $t$-structure, which can be defined as follows. The $v$-stack $\mathrm{Bun}_{G}$ is cohomologically smooth of $\ell$-dimension 00. Moreover, each one of the HN-strata $\mathrm{Bun}_{G}^{b}$ is isomorphic to $[\ast/\mathcal{J}_{b}]$, which is cohomologically smooth of $\ell$-dimension $-d_{b}=-\mathrm{dim}(\mathrm{Ig}^{b})$. Therefore, we can define a perverse $t$-structure $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (resp. $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$) on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ given by insisting that the !!! (resp. *) restrictions to $\mathrm{Bun}_{G}^{b}$ are concentrated in degrees $\geq\langle 2\rho_{G},\nu_{b}\rangle$ (resp. $\leq\langle 2\rho_{G},\nu_{b}\rangle$). The key result that follows from the work of [Ham2] and various compatibility results for the Fargues-Scholze correspondence is as follows.
@@
(Corollary 4.24) Let $\mu$ be a minuscule geometric dominant cocharacter and $G$ a product of groups satisfying the conditions of Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. Then if $\mathfrak{m}$ is generic the restriction of the Hecke operator
@@
is perverse $t$-exact. In particular, it induces maps
$j_{1}^{*}T_{\mu}:\phantom{}^{\mathrm{p}}\mathrm{D}^{\mathrm{ULA},\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\rightarrow\mathrm{D}^{\mathrm{adm},\geq 0}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}_{\phi_{\mathfrak{m}}}$
and
$j_{1}^{*}T_{\mu}:\phantom{}^{\mathrm{p}}\mathrm{D}^{\mathrm{ULA},\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\rightarrow\mathrm{D}^{\mathrm{adm},\leq 0}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})^{BW_{E_{\mu}}}_{\phi_{\mathfrak{m}}}$
on the halves of the perverse $t$-structure, where we note that the perverse $t$-structure on $\mathrm{D}(\mathrm{Bun}_{G}^{1},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ coincides with the usual $t$-structure.
Here is the table summarizing our local constraints:
[…]
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ $L=\mathbb{Q}_{p}$ $(\ell,2(p^{4}-1))=1$
$L/\mathbb{Q}_{p}$ unramified $(\ell,2[L:\mathbb{Q}_{p}](p^{4[L:\mathbb{Q}_{p}]}-1))=1$ $p\neq 2$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}])=1$
$G=\mathrm{U}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G=\mathrm{GU}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G(\mathrm{Sp}_{4,L})$ $L/\mathbb{Q}_{p}$ unramified, $L\neq\mathbb{Q}_{p}$ $(\ell,2[L:\mathbb{Q}_{p}](p^{4[L:\mathbb{Q}_{p}]}-1))=1$ $p\neq 2$
@@
Assuming the Fargues-Scholze correspondence for $G$ behaves as expected with rational coefficients, the analysis in [Ham2] allows one to verify this for any $\mu$ after imposing some additional conditions on the toral parameter $\phi_{T}^{\mathfrak{m}}$ attached to the maximal ideal $\mathfrak{m}$ ([Ham2, Condition/Definition 3.6]). However, for the groups considered, we will see that these additional conditions are superfluous and all one needs is generic, except for the case where $G=\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ or $G=G(\mathrm{Sp}_{4,L})$ with $L/\mathbb{Q}_{p}$ non-trivial, where we need an extra banality assumption on the prime $\ell$. It is conjectured ([Ham2, Conjecture 1.29]) that the results used to establish this theorem should always be true just under the condition that $\mathfrak{m}$ is generic.
These local torsion vanishing results would allow us to prove Conjecture 1.2 in several new cases if one could get control over the Igusa varieties $\mathrm{Ig}^{b}$. In Koshikawa’s argument, this is done by using a semi-perversity result proven by Caraiani-Scholze [CS2, Theorem 4.6.1], which was further generalized in work of Santos [San]. Roughly speaking, we want to show that $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ is concentrated in degrees $\geq d_{b}$, so that the complex of $J_{b}(\mathbb{Q}_{p})$-representations $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ defines the stalk of a semi-perverse sheaf on $\mathrm{Bun}_{G}$ at $b\in B(G)$, to which we can apply the previous result. In the case that the Shimura vareities $\mathrm{Sh}(\mathbf{G},X)_{K}$ are compact, there is a simpler way of seeing this. In particular, $\mathrm{Ig}^{b}$ is known to be a perfect affine scheme in this case, and so the desired semi-perversity just follows by applying Artin vanishing and then using Poincaré duality on the global Shimura variety. It turns out that this style of argument can be made to work even in the non-compact case. In [CS1, CS2, Ko, San], the non-compactly supported cohomology $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ is studied together with its filtration involving $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ coming from Mantovan’s formula, and shown to be concentrated in degrees $\geq d$. However, one could also study the compactly supported cohomology $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ and show that it is concentrated in degrees $\leq d$, à la Poincaré duality. To do this, we recall [CS2, Section 3.3] that, in the non-compact case, the perfect scheme $\mathrm{Ig}^{b}$ is not affine, but it admits a partial minimal compactification $g_{b}:\mathrm{Ig}^{b}\hookrightarrow\mathrm{Ig}^{b,*}$ which is affine, as proven in this more general setting of PEL type $A$ or $C$ by Santos [San]. We define
@@
the partially compactly supported cohomology, which is supported in degrees $\leq d_{b}$ by Artin-vanishing (Proposition 3.7). Now, for $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we define $\mathcal{S}(\mathbf{G},X)_{K}:=(\mathrm{Sh}(\mathbf{G},X)_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ to be the adic space over $\mathop{\rm Spa}(E_{\mathfrak{p}})$ attached to the Shimura variety. We can define the infinite level perfectoid Shimura varieties $\mathcal{S}(\mathbf{G},X)_{K^{p}}$ by taking the inverse limit of $\mathcal{S}(\mathbf{G},X)_{K^{p}K_{p}}$ as $K_{p}\rightarrow\{1\}$. The base-change $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is representable by a perfectoid space if $(\mathbf{G},X)$ is of pre-abelian type, and in general it is diamond. By the results of [Schtor, Han1], we have a Hodge-Tate period map
$\pi_{\mathrm{HT}}:[\mathcal{S}(\mathbf{G},X)_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$
recording the Hodge-Tate filtration on the abelian varieties with additional structure that $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ parametrizes. Here $\mathcal{F}\ell_{G,\mu^{-1}}:=(G_{C}/P_{\mu^{-1}})^{\mathrm{ad}}$ is the adic flag variety attached to the parabolic in $G_{C}$ given by a dominant inverse of $\mu$ and the dynamical method. We recall that the flag variety $[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ admits a locally closed stratification $i_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ indexed by $b\in B(G,\mu)$, given by pulling the HN-stratification along the natural map $h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G}$. We will now impose the following very mild assumption in what follows.
@@
Write $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$ for the closed complement of $\mathrm{Ig}^{b}$ in $\mathrm{Ig}^{b,*}$. We assume that $(\mathbf{G},X)$ is a PEL datum of type $A$ or $C$ such that, for all $b\in B(G,\mu)$, the perfect scheme $\partial\mathrm{Ig}^{b,*}$ is empty or has codimension in $\mathrm{Ig}^{b,*}$ greater than $2$.
@@
If $\mathbf{G}$ is simple then it is easy to show that this assumption will be satisfied if $\mathrm{dim}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})\geq 2$, by using that the boundary of the partially minimally compactified Igusa varieties is expressible as the Igusa varieties of Shimura varieties attached to Levis of $\mathbb{Q}$-rational parabolics of $\mathbf{G}$, as we will explain in §2.2.1. Moreover, if $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is compact then it is automatic that $\partial\mathrm{Ig}^{b,*}$ is empty. Therefore, if $\mathbf{G}$ is simple, this is excluding the cases where $\mathrm{dim}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})=1$ and $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is non-compact. There are two possibilities; either $(\mathbf{G},X)$ is the Shimura datum attached to the modular curve, or it is the Shimura datum attached to the unitary Shimura curve (See [Zha, Proposition 1.9]). In the latter case, we have that the connected components are given by modular curves. In these cases, the results of [Ko] are sufficient to prove Conjecture 1.2.
@@
Now, assuming this, one can show that the stalk of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at a geometric point $x:\mathop{\rm Spa}(C,C^{+})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}$ which lies in the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$ is given by $V_{b}$. Moreover, if we write $h_{b}^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\simeq\mathrm{Bun}_{G}^{b}$ for the pullback of $h^{\leftarrow}$ to $\mathrm{Bun}_{G}^{b}$ then one can deduce that the complex $i_{b!}i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ is isomorphic to $h^{\leftarrow*}j_{b!}(V_{b})$. Therefore, by excision, we deduce that the complex of $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-representations

### v5

@@ 1.0.2. Proof Sketch of the Main Theorem
for all possible $b\in B(G,\mu)_{\mathrm{un}}$. This makes the situation much more complicated; in fact, for non-split $G$, the basic element could be unramified, and in this case the Igusa variety $\mathrm{Ig}^{b}$ is just a profinite set, hence the problem of torsion vanishing for the contribution of the basic locus is completely reduced to controlling the generic part of the torsion cohomology of the local shtuka space attached to the basic element.
@@
Such control of the cohomology of shtuka spaces with torsion coefficients for these more general situations was attained in [24]. In order to understand this, it is helpful to move away from the language of isotypic parts of shtuka spaces and consider the action of Hecke operators on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$, the derived category of étale $\overline{\mathbb{F}}_{\ell}$-sheaves on $\mathrm{Bun}_{G}$. Since we are interested in cohomology localized at a generic maximal ideal $\mathfrak{m}$, Hansen constructs in appendix A a full-subcategory $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\subset\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ together with an idempotent localization map $(-)_{\phi_{\mathfrak{m}}}:\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})\rightarrow\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$ such that, on smooth irreducible representations, the localization map is either an isomorphism or 0 depending on if the representation has Fargues-Scholze parameter conjugate to $\phi_{\mathfrak{m}}$ or not (Lemma 4.3 (1)). We let $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ denote the full subcategory of ULA objects, where we recall by [16, Theorem V.7.1], that this is equivalent to insisting that the restrictions to all the HN-strata indexed by $b\in B(G)$ are valued in the full subcategories $\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ of admissible complexes (i.e the invariants under all open pro-$p$ $K\subset J_{b}(\mathbb{Q}_{p})$ is a perfect complex). Using the results of [24], we show (Corollary 4.29) that, under some regularity and genericity hypothesis on $\mathfrak{m}$, that one has a direct sum decomposition:
(2) $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}\simeq\bigoplus_{b\in B(G)_{\mathrm{un}}}\mathrm{D}^{\mathrm{adm}}(J_{b}(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}.$
@@
More precisely, we show that the !! and * push-forwards with respect to the inclusion of HN-strata agree on this sub-category, and so the excision semi-orthogonal decomposition splits on $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$. This decomposition is a refinement of the fact mentioned above that only the shtuka spaces corresponding to the unramified elements $b\in B(G,\mu)_{\mathrm{un}}$ can contribute to the generic localization of the cohomology of the Shimura variety. The desired control of the shtuka spaces is now in turn encoded in understanding how Hecke operators interact with a perverse $t$-structure on $\mathrm{Bun}_{G}$ after restricting to the localized category $\mathrm{D}^{\mathrm{ULA}}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi_{\mathfrak{m}}}$.
@@
$\mathrm{Hck}_{G,\leq\mu}$$\mathrm{Bun}_{G}$$\mathrm{Bun}_{G}\times\mathop{\rm Spd}(C),$$\scriptstyle{h_{\mu}^{\leftarrow}}$$\scriptstyle{h_{\mu}^{\rightarrow}}$
where $\mathrm{Hck}_{G,\leq\mu}$ is the stack parametrizing modifications $\mathcal{E}_{1}\dashrightarrow\mathcal{E}_{2}$ of a pair of $G$-bundles with meromorphy bounded by $\mu$ at the closed Cartier divisor defined by the fixed untilt given by $C$, and $h_{\mu}^{\rightarrow}$ (resp. $h_{\mu}^{\leftarrow}$) remembers $\mathcal{E}_{1}$ (resp. $\mathcal{E}_{2}$). We define
@@
[…]
@@
We are almost ready to state the result on Hecke operators we will need. To do this, we recall that $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ has a natural perverse $t$-structure, which can be defined as follows. The $v$-stack $\mathrm{Bun}_{G}$ is cohomologically smooth of $\ell$-dimension 0. Moreover, it has a locally closed HN-stratification $j_{b}:\mathrm{Bun}_{G}^{b}\hookrightarrow\mathrm{Bun}_{G}$, where each of the strata $\mathrm{Bun}_{G}^{b}$ are isomorphic to $\mathop{\rm Spd}(C)/\mathcal{J}_{b}]$, which is cohomologically smooth of $\ell$-dimension $d_{b}=-\mathrm{dim}(\mathrm{Ig}^{b})$. Therefore, we can define a perverse $t$-structure $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\geq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ (resp. $\phantom{}{}^{\mathrm{p}}\mathrm{D}^{\leq 0}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$) on $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})$ given by insisting that the !! (resp. *) restrictions to $\mathrm{Bun}_{G}^{b}$ are concentrated in degrees $\geq\langle 2\rho_{G},\nu_{b}\rangle$ (resp. $\leq\langle 2\rho_{G},\nu_{b}\rangle$). The key result that follows from the work of [24] and various compatibility results for the Fargues-Scholze correspondence is as follows.
@@
(Corollary 4.27) Let $\mu$ be a minuscule geometric dominant cocharacter and $G$ a product of groups satisfying the conditions of Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. Let $\mathfrak{m}$ be a generic maximal ideal with associated semi-simple $L$-parameter $\phi_{\mathfrak{m}}$. The restriction of the Hecke operator
@@
to the neutral strata $j_{1}:[\ast/\underline{G(\mathbb{Q}_{p})}]\simeq\mathrm{Bun}_{G}^{1}\hookrightarrow\mathrm{Bun}_{G}$ is perverse $t$-exact (in fact much more is true (See Corollary 4.27)), where we note that the perverse $t$-structure on $\mathrm{D}(\mathrm{Bun}_{G}^{1},\overline{\mathbb{F}}_{\ell})\simeq\mathrm{D}(G(\mathbb{Q}_{p}),\overline{\mathbb{F}}_{\ell})$ coincides with the usual $t$-structure.
@@
Assuming the Fargues-Scholze correspondence for $G$ behaves as expected with rational coefficients, the analysis in [24] allows one to verify this for any $\mu$ after imposing some additional conditions on the toral parameter $\phi_{T}^{\mathfrak{m}}$ attached to the maximal ideal $\mathfrak{m}$ such as being regular and satisfying Definition 4.16 for suitably chosen $\mu$, where the latter conditions are often implied by the genericity in the relevant use cases. We emphasize however that the above Theorem should always be true just under the condition that $\mathfrak{m}$ is generic, as formally stated in Conjecture 1.2.
These local torsion vanishing results would allow us to prove Conjecture 1.2 in several new cases if one could get control over the Igusa varieties $\mathrm{Ig}^{b}$. In Koshikawa’s argument, this is done by using a semi-perversity result proven by Caraiani-Scholze [8, Theorem 4.6.1], which was further generalized in work of Santos [54]. Roughly speaking, we want to show that $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ is concentrated in degrees $\geq d_{b}$, so that the complex of $J_{b}(\mathbb{Q}_{p})$-representations $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ defines the stalk of a semi-perverse sheaf on $\mathrm{Bun}_{G}$ at $b\in B(G)$, to which we can apply the previous result. In the case that the Shimura vareities $\mathrm{Sh}(\mathbf{G},X)_{K}$ are compact, there is a simpler way of seeing this. In particular, $\mathrm{Ig}^{b}$ is known to be a perfect affine scheme in this case, and so the desired semi-perversity just follows by applying Artin vanishing and then using Poincaré duality on the global Shimura variety. It turns out that this style of argument can be made to work even in the non-compact case. In [7, 8, 37, 54], the non-compactly supported cohomology $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ is studied together with its filtration involving $R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$ coming from Mantovan’s formula, and shown to be concentrated in degrees $\geq d$. However, one could also study the compactly supported cohomology $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ and show that it is concentrated in degrees $\leq d$, à la Poincaré duality. To do this, we recall [8, Section 3.3] that, in the non-compact case, the perfect scheme $\mathrm{Ig}^{b}$ is not affine, but it admits a partial minimal compactification $g_{b}:\mathrm{Ig}^{b}\hookrightarrow\mathrm{Ig}^{b,*}$ which is affine, as proven in this more general setting of PEL type AC by Santos [54]. We define
@@
the partially compactly supported cohomology, which is supported in degrees $\leq d_{b}$ by Artin-vanishing (Proposition 3.6). Now, for $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we define $\mathcal{S}(\mathbf{G},X)_{K}:=(\mathrm{Sh}(\mathbf{G},X)_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ to be the adic space over $\mathop{\rm Spa}(E_{\mathfrak{p}})$ attached to the Shimura variety. We can define the infinite level perfectoid Shimura varieties $\mathcal{S}(\mathbf{G},X)_{K^{p}}$ by taking the inverse limit of $\mathcal{S}(\mathbf{G},X)_{K^{p}K_{p}}$ as $K_{p}\rightarrow\{1\}$ in the category of diamonds. The base-change $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is representable by a perfectoid space if $(\mathbf{G},X)$ is of pre-abelian type, and in general it is diamond. By the results of [57, 28], we have a Hodge-Tate period map
$\pi_{\mathrm{HT}}:[\mathcal{S}(\mathbf{G},X)_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$
recording the Hodge-Tate filtration on the abelian varieties with additional structure that $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ parametrizes. Here $\mathcal{F}\ell_{G,\mu^{-1}}^{:}=(G_{C}/P_{\mu^{-1}})^{\mathrm{ad}}$ is the adic flag variety attached to the parabolic in $G_{C}$ given by a dominant inverse of $\mu$ and the dynamical method (i.e we look at the set $g\in G_{C}$ such that $\mathop{\rm lim}\nolimits_{t\rightarrow 0}\mathrm{ad}(\mu(t))g$ exists, see also Remark 3.13 for a discussion of the sign conventions we use) and $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ denotes the associated diamond, and we have passed to the associated $v$-stack quotients by the $G(\mathbb{Q}_{p})$-action. We recall that the flag variety $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ admits a locally closed stratification $i_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ indexed by $b\in B(G,\mu)$, given by pulling the HN-stratification along the natural map $h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G}$. We will now impose the following very mild assumption in what follows.
@@
Write $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$ for the closed complement of $\mathrm{Ig}^{b}$ in $\mathrm{Ig}^{b,*}$. We assume that $(\mathbf{G},X)$ is a PEL datum of type AC such that, for all $b\in B(G,\mu)$, the perfect scheme $\partial\mathrm{Ig}^{b,*}$ is empty or has codimension in $\mathrm{Ig}^{b,*}$ greater than or equal to $2$.
@@
If $\mathbf{G}^{\mathrm{ad}}$ is simple then it is easy to show that this assumption will be satisfied if $\mathrm{dim}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})\geq 2$, by using that the boundary of the partially minimally compactified Igusa varieties is expressible as the Igusa varieties of Shimura varieties attached to Levis of $\mathbb{Q}$-rational parabolics of $\mathbf{G}$, as we will explain in §2.2.2. Moreover, if $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is compact then it is automatic that $\partial\mathrm{Ig}^{b,*}$ is empty. Therefore, if $\mathbf{G}^{\mathrm{ad}}$ is simple, this is excluding the cases where $\mathrm{dim}(\mathcal{S}(\mathbf{G},X)_{K^{p},C})=1$ and $\mathcal{S}(\mathbf{G},X)_{K^{p},C}$ is non-compact. There are two possibilities; either $(\mathbf{G},X)$ is the Shimura datum attached to the modular curve, or it is the Shimura datum attached to the unitary Shimura curve (See [68, Proposition 1.9]). In the latter case, we have that the connected components are given by modular curves. In these cases, the results of [37] are sufficient to prove Conjecture 1.2.
@@
Now, assuming this, one can show that the stalk of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at a geometric point $x:\mathop{\rm Spa}(C,C^{+})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ which lies in the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ is given by $V_{b}$. Moreover, if we write $h_{b}^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\simeq\mathrm{Bun}_{G}^{b}$ for the pullback of $h^{\leftarrow}$ to $\mathrm{Bun}_{G}^{b}$ then one can deduce that the complex $i_{b!}i_{b}^{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ is isomorphic to $h^{\leftarrow*}j_{b!}(V_{b})$. Therefore, by excision, we deduce that the complex of $G(\mathbb{Q}_{p})\times W_{E_{\mathfrak{p}}}$-representations